OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method. Solve the equation for y to get a y-intercept of (0, -2) and a slope of -one third. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, -2). Let y equal zero and solve for x to find an x-intercept of (-

OpenStudy (anonymous):

@sai1234 @waterineyes

OpenStudy (anonymous):

3x + 9y = -18 3/-18x + 9/-18y = 1 x/-6+y/-2=1 x intercept=-6,0 y intercept=(0,-2)

OpenStudy (anonymous):

A. B. or C.

OpenStudy (anonymous):

option c

OpenStudy (anonymous):

Let x equal zero and solve for y to find a y-intercept of (0, -2). Let y equal zero and solve for x to find an x-intercept of (-